OpenStudy (anonymous):

[2.08] The formula for degrees Celsius is: C = 9 over 5 (F − 32), where F stands for degrees Fahrenheit Part 2: Determine how many degrees Fahrenheit 2 degrees Celsius is.

OpenStudy (shamim):

i know equation is c/5=(F-32)/9

OpenStudy (shamim):

F-32=9C/5

OpenStudy (anonymous):

yeah nw transpose 32 to RHS F=9C/5+32

OpenStudy (shamim):

F=9C/5+32

OpenStudy (shamim):

if C=2 F=?

OpenStudy (shamim):

F=9*2/5+32

OpenStudy (anonymous):

@supanick14 put 2 in place of c n u l get the ans